OpenStudy (maddiie):

Without graphing,determine whether the following pairs of lines are (a)parallel (b)perpendicular, or (c) neither parallel nor perpendicular 5x-6y=19 6x+5y=-30

OpenStudy (anonymous):

Solve for y y=mx+b Compare m: it determines parallel, perp

OpenStudy (anonymous):

5x-6y=19 -6y = -5x+19 y = (5/6)x-(19/6) Since this line is expressed in the standard linear form of y=mx+b where m is the slope, so the slope of this line = m1= 5/6 6x+5y=-30 5y = -6x+30 y = (-6/5)x+6 Similarily, since this line is expressed in the standard linear form of y=mx+b where m is the slope, so the slope of this line = m2= -6/5 Since m1*m2 = (5/6)*(-6/5) = -1,then these two lines are perpendicular to each other.